# Time-delayed Newton’s law of cooling with a finite-rate thermal quench. Impact on the Mpemba and Kovacs effects

### v1

@@ II Single quench
Since $1<\kappa_{0}<e\simeq 2.718<\tau^{-1}<\kappa_{1}$, the first term on the right-hand side of Eq. (14) dominates over the second in the regime $t\gg 1$. Nonetheless, the inclusion of the second term ensures that Eq. (14) provides an excellent approximation for $\mathcal{E}_{0}(t)$ for all $t$, especially when $t>2\tau$. The maximum deviation of the approximation given by Eq. (14) from $\mathcal{E}_{0}(t)$ occurs at $t=0$ and $\tau=\tau_{\max}=e^{-1}$, with a value of $\frac{8}{3}e^{-1}-1\simeq-0.019$.
@@
Suppose that two samples (A and B) undergo the quench described by Eq. (3), except that their initial bath temperatures differ, with $T_{\text{b,A}}^{-}>T_{\text{b,B}}^{-}$, while the quench time ($t=0$) and final bath temperature ($T_{\text{b}}^{-}$) remain the same for both. According to Eq. (8), their relative temperature difference evolves as
@@ III Double quench
Figure 3: Sketch of a generic double finite-rate quench. Here, $T_{\text{b}}^{-}=4$ (in arbitrary units), $T_{\text{b}}^{-*}=1$, $T_{\text{b}}^{+}=3$, $\sigma=0.1$, $\tau=0.36$, and (a) $t_{\text{w}}=0.2$, (b) $t_{\text{w}}=0.4$. The dashed lines represent the bath temperature $T_{\text{b}}(t)$, while the solid lines represent the system temperature $T(t)$. For comparison, also the cases of instantaneous quenches ($\sigma=0$) are included, where the dotted and dashed-dotted lines represent $T_{\text{b}}(t)$ and $T(t)$, respectively.
@@
Figure 4: Illustration of the protocols (a) A and (b) B described in Table 1. Here, $T_{\text{b}}^{\text{h}}=4$ (in arbitrary units), $T_{\text{b}}^{\text{c}}=1$, $\sigma=0.1$, $\tau=0.36$, and $t_{\text{w}}=0.4$. The dashed lines represent the bath temperature $T_{\text{b}}(t)$, while the solid lines represent the system temperature $T(t)$. For comparison, also the cases of instantaneous quenches ($\sigma=0$) are included, where the dotted and dashed-dotted lines represent $T_{\text{b}}(t)$ and $T(t)$, respectively.
@@ IV Impact of finite-rate quenches on the Mpemba effect
Figure 5: (a) Phase space for the apparent Mpemba effect in the absence of time delay ($\tau=0$). The lower and upper curves represent $\tilde{t}_{\text{w}}^{\min}$ and $\tilde{t}_{\text{w}}^{\max}$, respectively, as functions of the quench characteristic time $\sigma$. If $t_{\text{w}}<\tilde{t}_{\text{w}}^{\min}$, then $\Delta_{\sigma}^{0}(t)>0$ for $t\geq 0$; if, on the other hand, $t_{\text{w}}>\tilde{t}_{\text{w}}^{\max}$, then $\Delta_{\sigma}^{0}(t)<0$ for $t\geq 0$. Therefore, an apparent Mpemba effect occurs if and only if $\tilde{t}_{\text{w}}^{\min}<t_{\text{w}}<\tilde{t}_{\text{w}}^{\max}$ (shaded region). (b) Evolution of $T_{\text{A}}(t)$ (solid line), $T_{\text{B}}(t)$ (dashed-dotted line), $T_{\text{b,A}}(t)$ (dashed line), and $T_{\text{b,B}}(t)$ (dotted line) for the undelayed case ($\tau=0$) with $\sigma=0.1$, $t_{\text{w}}=0.75$, $T_{\text{b}}^{\text{h}}=4$ (in arbitrary units), and $T_{\text{b}}^{\text{c}}=1$.
@@ V Impact of finite-rate quenches on the Kovacs effect
Figure 9: Kovacs effect for $T_{\text{b}}^{\text{h}}=4$ (arbitrary units), $T_{\text{b}}^{\text{c}}=1$, $\tau=0.36$, and $t_{\text{w}}=0.3,0.6,0.9$, comparing instantaneous quenches ($\sigma=0$, dashed-dotted lines) and finite-rate quenches ($\sigma=0.1$, solid lines). (a) Bath temperature $T_{\text{b}}(t)$. (b) Sample temperature $T(t)$, where dotted lines represent $T(t)$ if the system were not quenched to $T(t_{\text{w}})$ at $t=t_{\text{w}}$.
Figure 9 compares the Kovacs effect for $\tau=0.36$ and three waiting times ($t_{\text{w}}=0.3,0.6,0.9$) under instantaneous ($\sigma=0$) and finite-rate ($\sigma=0.1$) quenches. Three key differences are observed: (i) the temperature slope $\dot{T}(t)$ is continuous at $t=0$ and $t=t_{\text{w}}$ for finite-rate quenches, whereas it is discontinuous for instantaneous quenches; (ii) the sample temperature $T_{\text{b}}^{\text{w}}$ at the waiting time is higher for finite-rate quenches than for instantaneous quenches; and (iii) the relative size of the Kovacs hump is deeper in the case of finite-rate quenches.
@@ VI Conclusions
The data that support the findings of this study are available from the author on reasonable request.
Acknowledgements.
This work was inspired by discussions during my participation in the long-term workshop “Frontiers in Non-equilibrium Physics 2024” (YITP-T-24-01). I sincerely appreciate the warm hospitality extended to me by Prof. Hayakawa during my stay at the Yukawa Institute for Theoretical Physics, Kyoto University, in July 2024. I acknowledge financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033.

### v2

# Time-delayed Newton’s law of cooling with a finite-rate thermal quench: Impact on the Mpemba and Kovacs effects

@@ II Single quench
Since $1<\kappa_{0}<e\simeq 2.718<\tau^{-1}<\kappa_{1}$, the first term on the right-hand side of Eq. (14) dominates over the second in the regime $t\gg 1$. Nonetheless, the inclusion of the second term ensures that Eq. (14) excellently approximates $\mathcal{E}_{0}(t)$ for all $t$, especially when $t>2\tau$. The maximum deviation of the approximation given by Eq. (14) from $\mathcal{E}_{0}(t)$ occurs at $t=0$ and $\tau=\tau_{\max}=e^{-1}$, with a value of $\frac{8}{3}e^{-1}-1\simeq-0.019$.
@@
Suppose that two samples (A and B) undergo the quench described by Eq. (3), except that their initial bath temperatures differ, with $T_{\text{b,A}}^{-}>T_{\text{b,B}}^{-}$, while the quench time ($t=0$) and final bath temperature ($T_{\text{b}}^{+}$) remain the same for both. According to Eq. (8), their relative temperature difference evolves as
@@ III Double quench
Figure 3: Sketch of a generic double finite-rate quench. Here, $T_{\text{b}}^{-}=4$ (in arbitrary units), $T_{\text{b}}^{-*}=1$, $T_{\text{b}}^{+}=3$, $\sigma=0.1$, $\tau=0.36$, and (a) $t_{\text{w}}=0.2$, (b) $t_{\text{w}}=0.4$. The dashed lines represent the bath temperature $T_{\text{b}}(t)$, while the solid lines represent the system temperature $T(t)$. For comparison, also the cases of instantaneous quenches ($\sigma=0$) are included, where the dotted and dash-dotted lines represent $T_{\text{b}}(t)$ and $T(t)$, respectively.
@@
Figure 4: Illustration of the protocols (a) A and (b) B described in Table 1. Here, $T_{\text{b}}^{\text{h}}=4$ (in arbitrary units), $T_{\text{b}}^{\text{c}}=1$, $\sigma=0.1$, $\tau=0.36$, and $t_{\text{w}}=0.4$. The dashed lines represent the bath temperature $T_{\text{b}}(t)$, while the solid lines represent the system temperature $T(t)$. For comparison, also the cases of instantaneous quenches ($\sigma=0$) are included, where the dotted and dash-dotted lines represent $T_{\text{b}}(t)$ and $T(t)$, respectively.
@@ IV Impact of finite-rate quenches on the Mpemba effect
Figure 5: (a) Phase space for the apparent Mpemba effect in the absence of time delay ($\tau=0$). The lower and upper curves represent $\tilde{t}_{\text{w}}^{\min}$ and $\tilde{t}_{\text{w}}^{\max}$, respectively, as functions of the quench characteristic time $\sigma$. If $t_{\text{w}}<\tilde{t}_{\text{w}}^{\min}$, then $\Delta_{\sigma}^{0}(t)>0$ for $t\geq 0$; if, on the other hand, $t_{\text{w}}>\tilde{t}_{\text{w}}^{\max}$, then $\Delta_{\sigma}^{0}(t)<0$ for $t\geq 0$. Therefore, an apparent Mpemba effect occurs if and only if $\tilde{t}_{\text{w}}^{\min}<t_{\text{w}}<\tilde{t}_{\text{w}}^{\max}$ (shaded region). (b) Evolution of $T_{\text{A}}(t)$ (solid line), $T_{\text{B}}(t)$ (dash-dotted line), $T_{\text{b,A}}(t)$ (dashed line), and $T_{\text{b,B}}(t)$ (dotted line) for the undelayed case ($\tau=0$) with $\sigma=0.1$, $t_{\text{w}}=0.75$, $T_{\text{b}}^{\text{h}}=4$ (in arbitrary units), and $T_{\text{b}}^{\text{c}}=1$.
@@ V Impact of finite-rate quenches on the Kovacs effect
Figure 9: Kovacs effect for $T_{\text{b}}^{\text{h}}=4$ (arbitrary units), $T_{\text{b}}^{\text{c}}=1$, $\tau=0.36$, and $t_{\text{w}}=0.3,0.6,0.9$, comparing instantaneous quenches ($\sigma=0$, dash-dotted lines) and finite-rate quenches ($\sigma=0.1$, solid lines). (a) Bath temperature $T_{\text{b}}(t)$. (b) Sample temperature $T(t)$, where dotted lines represent $T(t)$ if the system were not quenched to $T(t_{\text{w}})$ at $t=t_{\text{w}}$.
Figure 9 compares the Kovacs effect for $\tau=0.36$ and three waiting times ($t_{\text{w}}=0.3,0.6,0.9$) under instantaneous ($\sigma=0$) and finite-rate ($\sigma=0.1$) quenches. Three key differences are observed: (i) the temperature slope $\dot{T}(t)$ is continuous at $t=0$ and $t=t_{\text{w}}$ for finite-rate quenches, whereas it is discontinuous for instantaneous quenches; (ii) the sample temperature $T_{\text{b}}^{\text{w}}$ at the waiting time is higher for finite-rate quenches than for instantaneous quenches; and (iii) the Kovacs hump is deeper in the case of finite-rate quenches.
@@ VI Conclusions
Acknowledgements.
This work was inspired by discussions during my participation in the long-term workshop “Frontiers in Non-equilibrium Physics 2024” (YITP-T-24-01). I sincerely appreciate the warm hospitality extended to me by Prof. Hayakawa during my stay at the Yukawa Institute for Theoretical Physics, Kyoto University, in July 2024. I acknowledge financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe”.
@@ Acknowledgements.
The data that support the findings of this article are not publicly available because this publication is theoretical work. The data are available from the author upon reasonable request.